how to solve the non-linear differential equation implicitly.

y' = x + 2y / 2x + y

Please show your work step by step. Thanks.
Answer #1

Consider af ry da a 2xry not (2&ty Idy = (x+24) da (x+2y) dx - (2xty Jay =0 – Comparing with dee Ndy so weget 19-atay and H=-unde & f (terms in ext free from x) ay = e. Y-cont. e l x+24 dx of x2 - y2 4 - 6 -cont. Samaya de tallerne = 4-coust -6 Ś log